# Category $\mathcal{O}$ for Lie superalgebras

### v2

@@ Abstract.
The authors define a Category $\mathcal{O}$ for any quasi-reductive Lie superalgebra $\mathfrak{g}$ with respect to a triangular decomposition. This much needed approach unifies many important constructions in the existing literature in a rigorous fashion. Our Category $\mathcal{O}$ encompasses all highest weight categories for Lie (super)algebras as well as specific examples which may not be highest weight categories. When the decomposition arises from a principal parabolic subalgebra $\mathfrak{p}$ of $\mathfrak{g}$, the Category $\mathcal{O}$ exhibits rich homological properties. For one, the authors show that in contrast to the case of a semisimple Lie algebra, the Category $\mathcal{O}$ is standardly stratified. Moreover, the categorical cohomology of $\mathcal{O}$ is a finitely generated ring. This provides a first step towards developing a support variety theory for Category $\mathcal{O}$. Finally, it is shown that the complexity of modules in Category $\mathcal{O}$ is finite with an explicit upper bound given by the dimension of the subspace of the odd degree elements in $\mathfrak{g}$. This upgrades results known for $\mathfrak{gl}(m|n)$ to the more general setting. Our arguments are based on foundational connections between the categorical cohomology and the relative Lie superalgebra cohomology as well as the interplay between Category $\mathcal{O}$ for $\mathfrak{g}$ and the Category $\mathcal{O}$ for its corresponding Lie algebra $\mathfrak{g}_{\bar{0}}$.
@@ 3.1. The Definition
Let $\mathfrak{a}\subseteq\mathfrak{g}$ be a Lie sub-superalgebra and suppose that $\mathfrak{g}=\mathfrak{a}\oplus\mathfrak{a}^{\prime}$ as vector spaces for some vector space $\mathfrak{a}^{\prime}$.
@@
Let $\mathcal{C}_{(\mathfrak{g},\mathfrak{a})}$ be the full subcategory of $\mathfrak{g}$-Mod with the following properties.
  1. (C1)
    Every module $M$ is a direct sum of finite-dimensional $\mathfrak{a}$-modules.
  2. (C2)
    $\mathfrak{a}^{\prime}$ is stable under $\mathfrak{a}$.
@@
For the special case when $\mathfrak{g}=\mathfrak{g}_{\bar{0}}$ is a Lie algebra with $\mathfrak{a}=\mathfrak{t}_{\bar{0}}$ being its Cartan subalgebra, condition (C2) is equivalent to the usual weight module condition with respect to $\mathfrak{t}_{\bar{0}}$.
@@ 3.2. Principal Parabolic Subalgebras
Recall that a subalgebra of $\mathfrak{g}$ is parabolic if it contains a Borel subalgebra. Following [GY13], we consider those parabolic subalgebras which are  [^1] The authors thank Mazorchuk for informing us that such parabolic subalgebras appeared first in the work [DMP00] by Dimitrov, Mathieu, and Penkov.principal so that there is a triangular decomposition of $\mathfrak{g}$ in terms of a separating hyperplane. Let $\Phi$ be the set of roots of $\mathfrak{g}$ lying in a Euclidean space $V$.
@@
$\mathfrak{n}^{+}:=\bigoplus\nolimits_{\alpha\in N^{+}}\mathfrak{g}_{\alpha},\quad\mathfrak{n}^{-}:=\bigoplus\nolimits_{\alpha\in-(N^{+})}\mathfrak{g}_{\alpha},\quad\mathfrak{l}:=\mathfrak{t}\oplus\bigoplus\nolimits_{\alpha\in L}\mathfrak{g}_{\alpha}.$ (3.2.1)
@@ 3.3. Important facts about ${\mathcal{O}}^{{\mathfrak{p}}_{{\bar{0}}}}$
The category ${\mathcal{O}}^{{\mathfrak{p}}_{{\bar{0}}}}$ decomposes into direct summands ${\mathcal{O}}^{{\mathfrak{p}}_{{\bar{0}}}}_{\lambda}$ such that ${\mathcal{O}}^{{\mathfrak{p}}_{{\bar{0}}}}=\mathcal{\bigoplus}_{\lambda}{\mathcal{O}}^{{\mathfrak{p}}_{{\bar{0}}}}_{\lambda}$, where $\lambda$ runs over a complete set of representatives for the orbits under the dot action, and ${\mathcal{O}}^{{\mathfrak{p}}_{{\bar{0}}}}_{\mu}$ is the summand generated by the simple modules $L(\lambda)$ with $\lambda\in W_{{\mathfrak{p}}_{{\bar{0}}}}\cdot\mu$.
@@ 5.5. Cohomology for the pair $(\mathfrak{g},\mathfrak{h})$ and finite generation
(a) and (d) follow from [Ma19, Main Theorem]. (b) and (c) Let $M$ be a finite-dimensional $\mathfrak{g}$-module. Observe that $S(\mathfrak{g}_{{\bar{1}}}^{*})\otimes M$ is a finitely generated $S(\mathfrak{g}_{{\bar{1}}}^{*})$-module. Since $G_{{\bar{0}}}$ is reductive and the fixed points under $G_{{\bar{0}}}$ is exact, $C^{\bullet}(\mathfrak{g},\mathfrak{g}_{{\bar{0}}},M)$ is a finitely generated module over $S(\mathfrak{g}_{{\bar{1}}}^{*})^{G_{{\bar{0}}}}$. Therefore, by Theorem 5.3.1, $C^{\bullet}(\mathfrak{g},\mathfrak{g}_{{\bar{0}}},M)$ is a finitely generated module over $R=\operatorname{H}^{\bullet}(\mathfrak{g},\mathfrak{g}_{{\bar{0}}},{\mathbb{C}})$, thus any subquotient is finitely generated, and ${\operatorname{H}}^{\bullet}(\mathfrak{g},\mathfrak{g}_{{\bar{0}}},M)$ is finitely generated as an $R$-module.
@@ 6.2. Comparison between ${\mathcal{O}}^{\mathfrak{p}}$ and ${\mathcal{O}}^{{\mathfrak{p}}_{{\bar{0}}}}$
$\begin{split}(U({\mathfrak{g}})\otimes_{U({\mathfrak{g}}_{{\bar{0}}})}P_{{\bar{0}}}(\sigma):P(\lambda))&=\dim\text{Hom}_{{\mathcal{O}}^{\mathfrak{p}}}(U({\mathfrak{g}})\otimes_{U({\mathfrak{g}}_{{\bar{0}}})}P_{{\bar{0}}}(\sigma),L(\lambda))\\ &=\dim\text{Hom}_{U({\mathfrak{g}})}(U({\mathfrak{g}})\otimes_{U({\mathfrak{g}}_{{\bar{0}}})}P_{{\bar{0}}}(\sigma),L(\lambda))\\ &=\dim\text{Hom}_{U({\mathfrak{g}}_{{\bar{0}}})}(P_{{\bar{0}}}(\sigma),L(\lambda))\\ &=\dim\text{Hom}_{{\mathcal{O}}^{{\mathfrak{p}}_{{\bar{0}}}}}(P_{{\bar{0}}}(\sigma),L(\lambda))\\ &=[L(\lambda):L_{{\bar{0}}}(\sigma)].\end{split}$ (6.2.2)
@@ 6.4. Proof of Theorem A
$\begin{split}[R^{j}\mathcal{F}(N):L(\lambda)]&=\dim\mathrm{Ext}^{j}_{(\mathfrak{g},\mathfrak{l}_{\bar{0}})}(P(\lambda),N)\\ &\leq\dim\mathrm{Ext}^{j}_{(\mathfrak{g},\mathfrak{l}_{\bar{0}})}(U(\mathfrak{g})\otimes_{U(\mathfrak{g}_{\bar{0}})}P_{{\bar{0}}}(\sigma),N)\\ &=\dim\mathrm{Ext}^{j}_{(\mathfrak{g}_{{\bar{0}}},\mathfrak{l}_{\bar{0}})}(P_{{\bar{0}}}(\sigma),N)\\ &=\dim\mathrm{Ext}^{j}_{{\mathcal{O}}^{{\mathfrak{p}}_{{\bar{0}}}}}(P_{{\bar{0}}}(\sigma),N)\\ &=0.\end{split}$ (6.4.2)
Therefore, $R^{j}\mathcal{F}(N)=0$ for $j>0$.
@@
$\mathrm{Ext}^{n}_{(\mathfrak{g},\mathfrak{l}_{\bar{0}})}(M,N)\cong\mathrm{Ext}^{n}_{\mathcal{O}^{\mathfrak{p}}}(M,\mathcal{F}(N))\cong\mathrm{Ext}^{n}_{\mathcal{O}^{\mathfrak{p}}}(M,N).$ (6.4.4)
@@ 7.3. Proof of Theorem C
For Condition (d), we will construct a $\nabla$-filtration of $I(\lambda)$ or $\Pi I(\lambda$) using Theorem 7.2.2. Since $I(\lambda)$ is injective in ${\mathcal{O}}^{\mathfrak{p}}$, it follows that $\mathrm{Ext}^{n}_{\mathcal{O}^{\mathfrak{p}}}(\Delta_{\mathrm{irr}}(\gamma),I(\lambda))=0$ for all $\gamma\in\Lambda$ and $n>0$. Therefore, $I(\lambda)$ has a $\nabla_{\mathrm{inj}}$-filtration. The proof of Theorem 7.2.2 demonstrates that the filtration can be constructed with the necessary properties.
@@ 8.2. Universally $\bar{0}$}-bounded
The module $\Lambda^{\bullet}({\mathfrak{g}}_{{\bar{1}}})\otimes L_{{\mathfrak{p}}_{{\bar{0}}}}(\sigma)$ is a finite-dimensional module over $U({\mathfrak{l}}_{{\bar{0}}})$. One can show that the composition multiplicities are bounded by the number of weights of $\gamma=\sigma+\mu$ where $\mu$ is a weight of $\Lambda^{\bullet}({\mathfrak{g}}_{{\bar{1}}})$, and $\gamma$ is dominant with respect to ${\mathfrak{l}}_{{\bar{0}}}$. This number is clearly bounded by the dimension of $\Lambda^{\bullet}({\mathfrak{g}}_{{\bar{1}}})$. Now for each $\gamma$, the composition multiplicities of $Z_{{\bar{0}}}(\gamma):L_{{\bar{0}}}(\tau)]$ are bounded by Proposition 3.3.1(c). Hence, the composition factor multiplicities of the $U({\mathfrak{g}}_{{\bar{0}}})$-module in (8.2.2) is universally bounded for all $\sigma$, and thus the simple modules in ${\mathcal{O}}^{\mathfrak{p}}$ are universally $\bar{0}$}-bounded.
@@ 8.3. Proof of Theorem D
$c_{\mathfrak{p}}(M)\leq r(S^{\bullet}({\mathfrak{g}}_{{\bar{1}}}^{*}))=\dim{\mathfrak{g}}_{{\bar{1}}}.$ (8.3.7)

### v3

@@ Abstract.
The authors define a Category $\mathcal{O}$ for any quasi-reductive Lie superalgebra $\mathfrak{g}$ with respect to a triangular decomposition. This much needed approach unifies many important constructions in the existing literature in a rigorous fashion. Our Category $\mathcal{O}$ encompasses all highest weight categories for Lie (super)algebras as well as specific examples which may not be highest weight categories. When the decomposition arises from a principal parabolic subalgebra $\mathfrak{p}$ of $\mathfrak{g}$, the Category $\mathcal{O}$ exhibits rich homological properties. For one, the authors show that in contrast to the case of a semisimple Lie algebra, the Category $\mathcal{O}$ is standardly stratified.
Furthermore, the categorical cohomology of $\mathcal{O}$ is a finitely generated ring. This provides a first step towards developing a support variety theory for Category $\mathcal{O}$. It is shown that the complexity of modules in Category $\mathcal{O}$ is finite with an explicit upper bound given by the dimension of the subspace of the odd degree elements in $\mathfrak{g}$. This upgrades results known for $\mathfrak{gl}(m|n)$ to the more general setting. Our arguments are based on foundational connections between the categorical cohomology and the relative Lie superalgebra cohomology as well as the interplay between Category $\mathcal{O}$ for $\mathfrak{g}$ and the Category $\mathcal{O}$ for its corresponding Lie algebra $\mathfrak{g}_{\bar{0}}$.
@@ 3.1. The Definition
Let $\mathfrak{a}\subseteq\mathfrak{g}$ be a Lie sub-superalgebra. Suppose that $\mathfrak{g}=\mathfrak{a}\oplus\mathfrak{a}^{\prime}$ as vector spaces for some vector space $\mathfrak{a}^{\prime}$ which is stable under $\mathfrak{a}$.
@@
Let $\mathcal{C}_{(\mathfrak{g},\mathfrak{a})}$ be the full subcategory of $\mathfrak{g}$-Mod such that each object $M\in\mathcal{C}_{(\mathfrak{g},\mathfrak{a})}$ is a direct sum of finite-dimensional $\mathfrak{a}$-modules.
@@
For the special case when $\mathfrak{g}=\mathfrak{g}_{\bar{0}}$ is a Lie algebra with $\mathfrak{a}=\mathfrak{t}_{\bar{0}}$ being its Cartan subalgebra, $C_{(\mathfrak{g}_{0},\mathfrak{t}_{0})}$ is the category of weight modules in $\mathfrak{g}_{\bar{0}}$-Mod.
@@ 3.2. Principal Parabolic Subalgebras
Recall that a subalgebra of $\mathfrak{g}$ is parabolic if it contains a Borel subalgebra. Following [GY13], we consider those parabolic subalgebras which are  [^1] The authors thank Mazorchuk for informing us that such parabolic subalgebras appeared first in the work [DMP00] by Dimitrov, Mathieu, and Penkov.principal. Let $\Phi$ be the set of roots of $\mathfrak{g}$ lying in a Euclidean space $V$.
@@
$\mathfrak{n}^{+}:=\bigoplus_{\alpha\in N^{+}}\mathfrak{g}_{\alpha},\qquad\mathfrak{n}^{-}:=\bigoplus_{-\alpha\in N^{+}}\mathfrak{g}_{\alpha},\qquad\mathfrak{l}:=\mathfrak{t}\oplus\Big(\bigoplus_{\alpha\in L}\mathfrak{g}_{\alpha}\Big).$ (3.2.1)
@@ 3.3. Important facts about ${\mathcal{O}}^{{\mathfrak{p}}_{{\bar{0}}}}$
The category ${\mathcal{O}}^{{\mathfrak{p}}_{{\bar{0}}}}$ decomposes into direct summands ${\mathcal{O}}^{{\mathfrak{p}}_{{\bar{0}}}}_{\lambda}$ such that ${\mathcal{O}}^{{\mathfrak{p}}_{{\bar{0}}}}=\mathcal{\bigoplus}_{\lambda}{\mathcal{O}}^{{\mathfrak{p}}_{{\bar{0}}}}_{\lambda}$, where $\lambda$ runs over a complete set of representatives for the orbits under the dot action, and ${\mathcal{O}}^{{\mathfrak{p}}_{{\bar{0}}}}_{\mu}$ is the summand generated by the simple modules $L_{\bar{0}}(\lambda)$ with $\lambda\in W_{{\mathfrak{p}}_{{\bar{0}}}}\cdot\mu$. Let $Z_{\bar{0}}(\lambda)$ be the corresponding Verma module.
@@ 5.5. Cohomology for the pair $(\mathfrak{g},\mathfrak{h})$ and finite generation
(a) and (d) follow from [Ma19, Main Theorem]. (b) and (c) Let $M$ be a finite-dimensional $\mathfrak{g}$-module. Observe that $S(\mathfrak{g}_{{\bar{1}}}^{*})\otimes M$ is a finitely generated $S(\mathfrak{g}_{{\bar{1}}}^{*})$-module. Since $G_{{\bar{0}}}$ is reductive and the fixed points under $G_{{\bar{0}}}$ is exact, $C^{\bullet}(\mathfrak{g},\mathfrak{g}_{{\bar{0}}},M)$ is a finitely generated module over $S(\mathfrak{g}_{{\bar{1}}}^{*})^{G_{{\bar{0}}}}$. Therefore, by Theorem 5.3.1, $C^{\bullet}(\mathfrak{g},\mathfrak{g}_{{\bar{0}}},M)$ is a finitely generated module over $R=\operatorname{H}^{\bullet}(\mathfrak{g},\mathfrak{g}_{{\bar{0}}},{\mathbb{C}})$, thus any subquotient is finitely generated, and ${\operatorname{H}}^{\bullet}(\mathfrak{g},\mathfrak{g}_{{\bar{0}}},M)$ is finitely generated as an $R$-module. ∎
@@ 6.2. Comparison between ${\mathcal{O}}^{\mathfrak{p}}$ and ${\mathcal{O}}^{{\mathfrak{p}}_{{\bar{0}}}}$
$\begin{split}(U({\mathfrak{g}})\otimes_{U({\mathfrak{g}}_{{\bar{0}}})}P_{{\bar{0}}}(\sigma):P(\lambda))&=\dim\text{Hom}_{{\mathcal{O}}^{\mathfrak{p}}}(U({\mathfrak{g}})\otimes_{U({\mathfrak{g}}_{{\bar{0}}})}P_{{\bar{0}}}(\sigma),L(\lambda))\\ &=\dim\text{Hom}_{U({\mathfrak{g}})}(U({\mathfrak{g}})\otimes_{U({\mathfrak{g}}_{{\bar{0}}})}P_{{\bar{0}}}(\sigma),L(\lambda))\\ &=\dim\text{Hom}_{U({\mathfrak{g}}_{{\bar{0}}})}(P_{{\bar{0}}}(\sigma),L(\lambda))\\ &=\dim\text{Hom}_{{\mathcal{O}}^{{\mathfrak{p}}_{{\bar{0}}}}}(P_{{\bar{0}}}(\sigma),L(\lambda))\\ &=[L(\lambda):L_{{\bar{0}}}(\sigma)].\qed\end{split}$ (6.2.2)
@@ 6.4. Proof of Theorem A
$\begin{split}[R^{j}\mathcal{F}(N):L(\lambda)]&=\dim\mathrm{Ext}^{j}_{(\mathfrak{g},\mathfrak{l}_{\bar{0}})}(P(\lambda),N)\\ &\leq\dim\mathrm{Ext}^{j}_{(\mathfrak{g},\mathfrak{l}_{\bar{0}})}(U(\mathfrak{g})\otimes_{U(\mathfrak{g}_{\bar{0}})}P_{{\bar{0}}}(\sigma),N)\\ &=\dim\mathrm{Ext}^{j}_{(\mathfrak{g}_{{\bar{0}}},\mathfrak{l}_{\bar{0}})}(P_{{\bar{0}}}(\sigma),N)\\ &=\dim\mathrm{Ext}^{j}_{{\mathcal{O}}^{{\mathfrak{p}}_{{\bar{0}}}}}(P_{{\bar{0}}}(\sigma),N)=0.\end{split}$ (6.4.2)
Therefore, $R^{j}\mathcal{F}(N)=0$ for $j>0$. ∎
@@
$\mathrm{Ext}^{n}_{(\mathfrak{g},\mathfrak{l}_{\bar{0}})}(M,N)\cong\mathrm{Ext}^{n}_{\mathcal{O}^{\mathfrak{p}}}(M,\mathcal{F}(N))\cong\mathrm{Ext}^{n}_{\mathcal{O}^{\mathfrak{p}}}(M,N).\qed$ (6.4.4)
@@ 7.3. Proof of Theorem C
For Condition (d), we will construct a $\nabla$-filtration of $I(\lambda)$ or $\Pi I(\lambda$) using Theorem 7.2.2. Since $I(\lambda)$ is injective in ${\mathcal{O}}^{\mathfrak{p}}$, it follows that $\mathrm{Ext}^{n}_{\mathcal{O}^{\mathfrak{p}}}(\Delta_{\mathrm{irr}}(\gamma),I(\lambda))=0$ for all $\gamma\in\Lambda$ and $n>0$. Therefore, $I(\lambda)$ has a $\nabla_{\mathrm{inj}}$-filtration. The proof of Theorem 7.2.2 demonstrates that the filtration can be constructed with the necessary properties. ∎
@@ 8.2. Universally $\bar{0}$}-bounded
The module $\Lambda^{\bullet}({\mathfrak{g}}_{{\bar{1}}})\otimes L_{{\mathfrak{p}}_{{\bar{0}}}}(\sigma)$ is a finite-dimensional module over $U({\mathfrak{l}}_{{\bar{0}}})$. One can show that the composition multiplicities are bounded by the number of weights of $\gamma=\sigma+\mu$ where $\mu$ is a weight of $\Lambda^{\bullet}({\mathfrak{g}}_{{\bar{1}}})$, and $\gamma$ is dominant with respect to ${\mathfrak{l}}_{{\bar{0}}}$. This number is clearly bounded by the dimension of $\Lambda^{\bullet}({\mathfrak{g}}_{{\bar{1}}})$. Now for each $\gamma$, the composition multiplicities of $Z_{{\bar{0}}}(\gamma):L_{{\bar{0}}}(\tau)]$ are bounded by Proposition 3.3.1(c). Hence, the composition factor multiplicities of the $U({\mathfrak{g}}_{{\bar{0}}})$-module in (8.2.2) is universally bounded for all $\sigma$, and thus the simple modules in ${\mathcal{O}}^{\mathfrak{p}}$ are universally $\bar{0}$}-bounded. ∎
@@ 8.3. Proof of Theorem D
$c_{\mathfrak{p}}(M)\leq r(S^{\bullet}({\mathfrak{g}}_{{\bar{1}}}^{*}))=\dim{\mathfrak{g}}_{{\bar{1}}}.\qed$ (8.3.7)